Biting an unpopped kernel of popcorn hurts! As an experiment, a self-confessed connoisseur of cheap popcorn carefully counted 773 kernels and put them in a popper. After popping, the unpopped kernels were counted. There were 86. a) Construct a 90 percent confidence interval for the proportion of all kernels that would not pop. (b) Check the normality assumption. (c) Try the Very Quick Rule. Does it work well here? Why, or why not? (d) Why might this sample not be typical?: Biting an unpopped kernel of popcorn hurts! As an experiment, a self-confessed connoisseur of cheap popcorn carefully counted 773 kernels and put them in a popper. After popping, the unpopped kernels were counted. There were 86. a) Construct a 90 percent confidence interval for the proportion of all kernels that would not pop. (b) Check the normality assumption. (c) Try the Very Quick Rule. Does it work well here? Why, or why not? (d) Why might this sample not be typical?
Answer by stanbon(18999)

Biting an unpopped kernel of popcorn hurts! As an experiment, a self-confessed connoisseur of cheap popcorn carefully counted 773 kernels and put them in a popper. After popping, the unpopped kernels were counted. There were 86.
------------------
p-hat = 86/773 = 0.1113
E = 1.645*sqrt[(0.1113)(0.8887)/773) = 0.0164..
-----------------------
a) Construct a 90 percent confidence interval for the proportion of all kernels that would not pop.
(0.1113-0.0164 < p < 0.1113+0.0164)
----------------------------------------
(b) Check the normality assumption.
I'll leave that to you
-----------------------------
(c) Try the Very Quick Rule. Does it work well here? Why, or why not?
I don't know what your Quick Rule might be.
-----------------------------
(d) Why might this sample not be typical?
It is based on only one sample of popcorn.
=========================================
cheers,
Stan H.